# Talking Wikidata: Communication patterns and their impact on community engagement in collaborative knowledge graphs

### v1

@@ 5.1 RQ1 - What are the characteristics of editors’ collaborations?
We conducted network analysis and descriptive statistics. We built and analysed three editor networks: (i) with all editors found in the discussions dataset (NF); (ii) with human editors, registered and unregistered (NH); and (iii) with only registered editors (NR). In each network, we considered nodes as editors, humans or bots, who had at least one post on a Wikidata discussion page, and edges were formed between editors participating in the same thread.
@@
Wikidata bots, like Wikipedia bots, participated in discussions either through bot controllers or by delivering automatic messages [28]. We elaborate more in Section 6.1.1.
@@
We employed centrality metrics (see Table 3) [5] to investigate and compare the three editor networks. We then further explored editors’ behaviour in the NR network. Specifically, we investigated whether NR forms a small world network by analysing the connected components in the network (which is a connected sub-network that is not part of any other larger network connected component) and measuring the average clustering coefficient and average shortest path. A small world network forms a giant component and presents a high average clustering coefficient and low average shortest path [52]. The clustering coefficient (as described in Table 3) measures the degree to which nodes in a network cluster together. It quantifies the likelihood that two neighbours of a node are also connected to each other. The shortest path refers to the minimum number of steps or edges required to traverse from one node to another node in a network. To test our small world hypothesis, we compared these metrics with a similar small world network generated with the Watts-Strogatz method [52], as well as with a similar random network. We then explored network characteristics like the distribution of clustering coefficient, shortest paths, and assortativity [32].
@@
Finally, we analysed the characteristics of registered editors using the features listed in Tables 3.
@@
5.2 RQ2 - What impacts whether a discussion will receive a response?
We developed a hybrid machine learning model to study the factors that influence discussions in Wikidata. The aim was to predict the likelihood of a response to a post based on various features. The features included information for the editors who wrote the post, graph embedding that captured the topology of the discussion network as depicted in Figure 4, and text embedding that represented the content of the post.
@@ 5.2 RQ2 - What impacts whether a discussion will receive a response?
For the set of editors’ features (EF), we used account age (EFage), # of edits (EFedits), # of posts (EFposts), and access level (EFaccess level) from Table 3.
For graph embeddings (G), we created a discussion network including three types of nodes: editors, posts, and thread titles. Editors were connected to their posts, while posts were connected to the threads they belonged to (see Figure 4). To create the graph embedding vectors for this network, we experimented with several state-of-the-art models such as RotatE [42], TransR [26], and CompGCN [47]. We chose CompGCN with embedding dimensions equal to $100$ to form a graph embedding vector for each entity in this network. Due to the network structure and the low number of relations, CompGCN showed the best performance with Hits@1 $0.37$, Hits@3: $0.56$, Hits@5: $0.64$, Hits@10: $0.75$ and Mean Reciprocal Rank $0.5$.
To answer $RQ2$ we used the graph embedding vectors for the posts (Gpost).
For text representations (T) we used the $text-embedding-ada-002$, a $GPT3.5$ large language model from OpenAI. This model was capable of processing long documents like post discussions effectively [33]. We generated two types of text embeddings: one for each post (Tpost) and one for the entire thread from the first post to the ith post (Tthread). Prior to generating these embeddings, we preprocessed the raw data by removing signatures and timestamps at the end of the post, replacing the item and property alphanumeric codes (Q# and P#) with [ITEM] and [PROP], respectively, and any code elements with [URL] or [CODE]. Finally, we removed any remaining unnecessary punctuation.
To develop the hybrid machine learning model, we used a neural network and conducted an ablation study, which means testing each of the model features individually as well as in combinations. To train the model, we split our dataset to train, test, and validate sets of $60\%$, $20\%$, and $20\%$, respectively. We used a Multi-Layer Perceptron (MLP) with 13131\-313 layers (test combinations of $8$, $32$, and $64$ neurons (Section 6 presents the best results), with “relu” as activation function. The output layer consisted of a single neuron with “sigmoid” activation function. For training the network, we used “adam” as the optimizer with a learning rate set at $0.001$ and the “binary_crossentopy” loss. Furthermore, we used early stopping to stop the training if the validation loss did not improve in $100$ epochs. We evaluated results using precision, recall, accuracy, and F1.
5.3 RQ3 - Do discussions impact the editors’ engagement?
To address the $RQ3$, we emplyied a similar approach as in $RQ2$. This time, we delved into the factors that affect editors’ engagement, which we defined using the feature activity status as described above (see Table 3). This feature indicated whether an editor was active or inactive.
As in $RQ2$, we leveraged various features for editors, including their graph embedding vectors (Geditors), and text embedding vectors for their posts. For the latter, we estimated the average of all post text embeddings of each editor (Tavr). We then combined all these features to create a hybrid machine learning model.
@@ 5.3 RQ3 - Do discussions impact the editors’ engagement?
5.4 Sample dataset
The full discussion network, as depicted in Figure 4, was composed of more than $2.5M$ triples, which required a high computational power to estimate the graph embedding vectors. Therefore, we used a smaller sample of our data to answer $RQ2$ and $RQ3$ questions. Specifically, we analysed threads found in item talk pages, as they presented the best balance between posts receiving a response and those not ($44\%$ of posts had a response, $56\%$ did not). In addition, our previous study on Wikidata discussions showed that in item talk pages, only $50\%$ of posts starting a new thread receive community attention to get an answer [22]. This study aims to investigate further and reveal factors that influence responses and improve communication within the project.
For $RQ2$, we analysed a sample of $44,139$ posts, while for $RQ3$, we studied a sample of $6,596$ registered editors. In both cases, the graph embeddings were created based on the discussion network shown in Figure 4 with $75,055$ nodes ( $6,596$ editors, $44,139$ posts, $24,320$ threads) and $88,092$ edges/triples.
@@ 6.1.1 Comparison of the three editor networks
Figure 5 presents centrality metrics for the NF, NH, and NR networks. For all networks, degree centrality, which measures “how many friends I have”, presented a power law distribution with a high number of editors with no edges in all networks: $14\%$ NF, $24\%$ NH, $26\%$ NR. This indicated that a quarter of human editors did not receive responses when starting a discussion. Furthermore, $47\%$ of them are inactive, meaning that half of the editors who did not receive a response drop out, while the rest continued contributing without discussing. To further investigate this relation, we analysed whether the activity status of an editor has a relation with whether their posts received responses using the Chi-square test for categorical variables [46]. We found a p-value less than 0.05, which showed that these two features are related.
For eigenvector and clustering coefficient metrics, the three networks presented similar distributions (see Figure 5). The eigenvector shows “how well-connected are my friends” and the clustering coefficient shows “if a friend of a friend is my friend”. We did not observe any nodes with superior or inferior connected friends or clustering for registered, unregistered, and bots.
In terms of closeness, which measures “the distance I have with my friends”, we observed that NF nodes have shorter distances (high closeness), while NH and NR showed similar distribution. At the same time, betweenness, which indicates “how many times I connect friends”, presented lower values for NF. The higher closeness values and lower betweenness values for NF suggested the significance of bots in the network. Bots were the most connected editors in the network and helped in bridging distances between nodes, thereby making paths for all nodes shorter.
After the analysis, we looked into the nodes with the highest centrality metrics for the three networks. In the NF network, a bot had the highest degree, closeness, and betweenness values. “DeltaBot” was the most active editor in discussions. It responded to requests for deletion or merging of items with prefixed answers such as “{{on hold}} This item is linked from 3 others” after checking their status. For NH and NR networks, editors with administrative access levels, such as property creators or administrators, received the highest metrics. The analyses indicated the importance of bots in communication related to maintenance jobs, and the dedication of the editors with higher level access to communication.
@@ 6.1.2 Investigation of registered editor network
We conducted a study to verify the hypothesis that Wikidata editors’ communication interactions form a small world network. We first examined the connected components for the NR. We discovered that there are $5,722$ connected components, with the largest component containing $73\%$ of the nodes, which amounts to $16,169$ nodes and $470,126$ edges. The second-largest component included only $0.04\%$ of the nodes. This indicated that NR has a giant component and many very small components.
Next, we measured and analysed the average cluster coefficient and the average shortest path of NR. We found that the average clustering coefficient was $0.6$ and the average shortest path was $3$. In comparison, a similar small world network had similar characteristics with an average clustering coefficient of $0.1$ and an average shortest path of $3$. In contrast, a random network with similar characteristics had a much lower average clustering coefficient of only $0.004$ and an average shortest path of $3$. These comparisons, combined with the giant component, indicated that NR forms a small world network.
We then measured degree assortativity to investigate the tendency of editors to connect with similar to them editors. We found that NR had assortativity at $-0.2$. This leans to a disassortative network. This means that editors tent to talk with dissimilar to them others, e.g., editors with many connections tent to talk with editors with few connections.
@@
Figure 7 presents the clustering coefficient versus the degree of a node in the NR. The figure showed that a high number of nodes had both high clustering coefficient and degree. Such nodes were likely to be influential or central within their respective clusters or communities and lead to robust networks with multiple paths for information or interactions to flow. As a result, the network was resilient to the removal of individual nodes. It is worth noting that this could be the reason that NF, NH and NR presented similar clustering coefficient distributions in Section 6.1.1. Despite bots being the most connected editors in NF, their removal in NH and NR did not influence node connections.
We also conducted a study on the distribution of the shortest paths in the NR. Similar to the MSN study [25], we randomly sampled $1000$ nodes and calculated their shortest paths to all other nodes. Figure 7 shows the distribution of shortest paths. We observed that the longest shortest path is seven, which indicates that long paths did not exist in NR. By examining the percentile of shortest paths, we discovered that one can reach $90\%$ of nodes in just four steps.
6.1.3 Investigation of registered editors characteristics.
@@
Figure 8 illustrates the distribution of registered editors based on the number of edits, posts and account age. Our analysis revealed that both the number of edits and posts followed a power law distribution. The data showed that approximately $80\%$ of editors had less than $5$K edits and ten posts, indicating that they contributed very little and did not engage in discussions. This implied that only a small percentage of editors ($20\%$) were actively participating in discussions, suggesting that they were the ones making the key decisions within the Wikidata community.
Furthermore, we analysed the distribution of account age and found that $24\%$ of editors had stayed in the community for less than a year, and $10\%$ for editors had remained active for more than ten years. This trend has been consistent with the first four years of Wikidata [38].
We further conducted an investigation to determine which editors made high contributions to discussions based on their access level and account age. Figure 9 shows the number of posts made by different access roles to account age. The figure showed that editors who had been active on Wikidata for over six years were the most active in discussions. Property creators were the most active, followed by basic level users.
@@ 6.1.4 Summary
Furthermore, we found that only $20\%$ of editors were actively involved in discussions and decision-making, with most of them being members for more than six years. It is interesting to note that the majority of editors who participated in discussions were basic level users, and some of them had the highest number of posts.
6.2 RQ2 - What impacts whether a discussion will receive a response?
The study aimed to investigate the factors that influence whether a post on Wikidata could receive a response. The ablation study included text embedding for the post content, graph embedding for the post topology (Gpost) (see Figure 4), and features for the editor who wrote the post. We considered the editors’ features (EF), account age (EFage), number of edits (EFedits), number of posts (EFposts), and access level (EFaccess level). We also experimented with different text embeddings. We tested whether the use of text embeddings for only the first posts of threads (TfisrtPost) performed better than all posts in threads (Tpost), or the entire thread from the first to the ith post (Tthread).
The results of the experiment are presented in Table 4, which demonstrates that, between individual features, graph embedding performed the best in predicting whether a post would receive a response. This result was expected since a post that was part of a highly connected thread, as shown in Figure 4, had a higher chance of receiving a response.
In addition, we found high performance for the text embeddings with minimal variations between the TfirstPost, Tpost, and Tthread. This suggested that the content of the text played a significant role in eliciting a response. This was particularly important for TfirstPost results, which initiated a discussion. The content of a post affected whether or not the conversation continues with a response to the thread.
The best features for predicting whether a post would receive a response were its content and where it was posted (Gpost + TfirstPost). The characteristics of the editor, such as their account age and access level, did not seem to significantly impact whether the post would receive a response. This suggests that it was more important to focus on the content of the post and where it was placed rather than who wrote it.
Table 4: Results for the model predicting whether a post will receive a response.

### v2

@@ 5.1 RQ1 - What are the characteristics of editors’ collaborations?
We conducted network analysis and descriptive statistics. Network analysis is used across various fields due to its ability to reveal complex relationships and patterns within data. Its ability to reveal hidden patterns, optimize network performance, and identify key nodes makes it an invaluable tool across a wide range of disciplines. We built and analysed three editor networks: (i) with all editors found in the discussions dataset (NF); (ii) with human editors, registered and unregistered (NH); and (iii) with only registered editors (NR). In each network, we considered nodes as editors (humans or bots) who had at least one post on a Wikidata discussion page. We added edges between editors participating in the same thread.
@@
Wikidata bots, like Wikipedia bots, participated in discussions either through bot controllers or by delivering automatic messages [28]. We elaborate in Section 6.1.1.
We employed centrality metrics (see Table 3) [5] to investigate and compare the three editor networks. Centrality measures help to explore the characteristics of the Wikidata network of editors. They identify the most important or influential nodes in the network and explain the connection of nodes. This can be critical for targeting interventions or understanding power dynamics.
@@
We then further explored the behaviour of editors in the NR network. Specifically, we investigated whether NR forms a small world network by analysing the connected components in the network (which is a connected sub-network that is not part of any other larger network connected component) and measuring the average clustering coefficient and average shortest path. This analysis helps to explore the communication characteristics of the registered editors. It examines the robustness of the network and the connectivity of groups. A small world network forms a giant component and presents a high average clustering coefficient and low average shortest path [52]. The clustering coefficient (as described in Table 3) measures the degree to which nodes in a network cluster together. It quantifies the likelihood that two neighbours of a node are also connected to each other. The shortest path refers to the minimum number of steps or edges required to traverse from one node to another node in a network. To test our small world hypothesis, we compared these metrics with a similar small world network generated with the Watts-Strogatz method [52], as well as with a similar random network.
We then explored network characteristics like the distribution of clustering coefficient, shortest paths, and assortativity [32]. This helps to understand the structural tendencies of the network, indicating whether it is more segmented into active subgroups or more integratively connected.
@@
Finally, we analysed the characteristics of registered editors using the features listed in Tables 3. This analysis can provide insights into the structural properties of the network of editors to explore the behaviour of editors and engagement patterns.
@@
5.2 RQ2 - What factors affect whether a discussion will receive a response?
We developed a machine learning model to study the factors that affect discussions in Wikidata. The aim was to predict the likelihood of a response to a post based on various features. We combined features of: characteristics of the editors who wrote the post; graph embedding which are a type of representation learning where a graph is converted into a low dimensional space in which the graph information is preserved [cai2018comprehensive]; and text embedding, which are numerical representations of text that capture the semantic meaning of words, phrases, sentences, or entire documents [33]. The list of features used for the ML model is presented in Table 4. We conducted an ablation study, which means systematically removing or masking individual features (or groups of features) and observing the change in model performance. Then, we compared the performance metrics (e.g., accuracy and precision) of the model before and after removing each feature. This analysis helped us to understand which factors, either individually or in combination, affect the prediction of the posts.
Table 4: The list of features used in the ML model with descriptions.
Feature Description
EFage Account age: The range in months between the editor’s registration date and their last edit timestamp.
EFedits # of edits: The total number of edits excluding discussion page edits.
EFposts # of posts: The total number of posts in discussion pages.
EFaccess level Access level: The role/roles of editors based on Wikidata hierarchy.
EF All the above features.
Gpost Graph embeddings for posts.
Geditors Graph embeddings for editors.
TfirstPost Text embeddings for the first posts of each thread.
Tpost Text embeddings for individual posts
Tthread Text embeddings for each post combining the previous posts in the thread. This is the average of text embeddings for the first to the ith post
Tavr Text embeddings for the editors. This is the average of the text embeddings for all their posts.
For the features of editors (EF), we used the set of features (Table 3) proposed in Section 5.1 for the investigation of the characteristics of editor’s collaboration. We used account age (EFage), number of edits (EFedits), number of posts (EFposts), and access level (EFaccess level).
For graph embeddings (G), we created a discussion network including three types of nodes: editors, posts, and thread titles. Editors were connected to their posts, while posts were connected to the threads they belonged (Figure 4). The graph included $75,055$ nodes ($6,596$ editors, $44,139$ posts, $24,320$ threads) and $88,092$ edges. To create the graph embedding vectors for this network, we experimented with several types of state-of-the-art models. We experimented with translational distance models like TransR [26], RotatE [42], which employs rotations in the complex vector space, and the CompGCN [47], which combines the principles of graph convolutional networks with compositional operators. Due to the network structure and the low number of relations, none of the tested models presented good performance except for CompGCN. For CompGCN we applied embedding dimensions equal to $100$ to form a graph embedding vector for each entity in this network (i.e., editors, posts, and threads). CompGCN showed the best performance with Hits@1 $0.37$, Hits@3: $0.56$, Hits@5: $0.64$, Hits@10: $0.75$ and Mean Reciprocal Rank $0.5$.
The full discussion network was composed of more than $2.5M$ edges, which required high computational power to estimate the graph embedding vectors. Therefore, we used only discussions found in item talk pages as a sample and not the whole discussion dataset as described in Section 4. We chose this sample because item talk pages presented the best balance between posts receiving a response and those not ($44\%$ of posts had a response, $56\%$ did not). Our previous study on Wikidata discussions also showed that in item talk pages, only $50\%$ of posts starting a new thread receive community attention to get an answer [22]. In addition, as shown in Koutsiana et al. [22], item talk pages include a wide range of topics and issues, such as KE topics, fact accuracy, conflicts, and policies. These topics can reflect better than other discussion channels, such as property talk pages and project chat, the conversations across Wikidata [22].
To answer $RQ2$ we used the graph embedding vectors generated for the posts (Gpost).
@@ 5.2 RQ2 - What factors affect whether a discussion will receive a response?
For text representations (T) we used the $text-embedding-ada-002$, a $GPT3.5$ large language model from OpenAI. We used this model because it was among the best performing models in the literature, and it was capable of processing long documents like post discussions effectively [33]. We generated two types of text embeddings: one for each post (Tpost) and one for the entire thread from the first to the ith post (Tthread). Prior to generating these embeddings, we preprocessed the raw data by removing signatures and timestamps at the end of the post, replacing the item and property alphanumeric codes (Q# and P#) with [ITEM] and [PROP], respectively, and any code elements with [URL] or [CODE]. Finally, we removed any remaining unnecessary punctuation.
To develop the machine learning model, we used a neural network and concatenated the above features. Figure 5 illustrates the steps of the ML model. We used a Multi-Layer Perceptron (MLP) because it is a versatile, efficient, and straightforward choice for a wide range of machine learning tasks, particularly when dealing with structured data and simpler relationships [khalil2009performance]. To train the model, we randomly split our dataset to train, test, and validate sets of $60\%$, $20\%$, and $20\%$, respectively. The MLP had 13131\-313 layers (test combinations of $8$, $32$, and $64$ neurons) with “relu” as activation function. The output layer consisted of a single neuron with “sigmoid” as an activation function. For training the network, we used “adam” as the optimizer with a learning rate set at $0.001$ and the “binary_crossentopy” loss. Furthermore, we used early stopping to stop the training if the validation loss did not improve in $100$ epochs. We evaluated results using precision, recall, accuracy, and F1.
Figure 5: A diagram with the steps for the ML model.
5.3 RQ3 - Do discussions affect the editors’ engagement?
To address the $RQ3$, we employed the same approach as in $RQ2$. We used the same ML model (Figure 5). This time, we delved into the factors that affected the engagement of editors, which we defined using the feature activity status as described above (see Table 3). This feature indicated whether an editor was active or inactive.
As in $RQ2$, we leveraged various features for editors Table 4), including their graph embedding vectors (Geditors), and text embedding vectors for their posts. For the latter, we estimated the average of text embeddings for all posts of each editor (Tavr).We then concatenated all these features to the ML model (Figure 5). We evaluated results using precision, recall, accuracy, and F1.
@@ 6.1.1 Comparison of the three editor networks
Figure 6 presents centrality metrics for the NF, NH, and NR networks. For all networks, degree centrality, which measures the number of edges for a node, presented a power law distribution with a high number of editors with no edges in all networks: $14\%$ NF, $24\%$ NH, $26\%$ NR. This indicated that a quarter of human editors did not receive responses when starting a discussion. Furthermore, $47\%$ of them were inactive, meaning that half of the editors who did not receive a response drop out, while the rest continued contributing without discussing. In contrast, $68\%$ of editors who received a response are still active. To further investigate this relation, we analysed whether the activity status of an editor was related to their posts received responses using the Chi-square test for categorical variables [46]. We found a p-value less than 0.05, which showed that these two features were related.
For eigenvector and clustering coefficient metrics, the three networks presented similar distributions (Figure 6). The eigenvector shows the centrallity of a node as proportional to its neighbours’ importance, and the clustering coefficient quanitfies how close its neighbours are to a complete graph. We did not observe any nodes with superior or inferior connections or clustering for registered editors, unregistered editors, and bots. This means that the removal of bots or unregistered editors did not significantly affect the network’s overall structure and influence distribution. The network might have a high degree of connectivity and robustness, where the influence and connections are well-distributed among many nodes. This makes the network resilient to the removal of nodes. We explore this further below in Section 6.1.2.
In terms of closeness, which counts the average of shortest paths, we observed that NF nodes had shorter distances (high closeness), while NH and NR showed similar distribution. At the same time, betweenness, which counts the number of times a node acts like a bridge, presented lower values for NF. The higher closeness values and lower betweenness values for NF suggested that bots were the most connected editors in the network and helped in bridging distances between nodes, thereby making paths for all nodes shorter. However, the eigenvector and clustering coefficient showed that even when removing bots, the network stays resilient.
After the analysis, we looked into the nodes with the highest centrality metrics for the three networks. In the NF network, a bot had the highest degree, closeness, and betweenness values. “DeltaBot” was the most active editor in discussions. It responded to requests for deletion or merging of items with prefixed answers such as “{{on hold}} This item is linked from 3 others” after checking their status. However, as described in Section 2, the role of bots is to perform repeated tasks, and the purpose of this communication is to respond to requests with a single post. Bots do not interact with editors or participate in conversations. For NH and NR networks, editors with administrative access levels, such as property creators or administrators, received the highest metrics. The analyses indicated the importance of bots in communication related to maintenance jobs, and the dedication of the editors with higher level access to communication.
@@ 6.1.2 Investigation of registered editor network
We checked whether the communication interactions of Wikidata editors formed a small world network. We first examined the connected components for the NR network. We discovered that there were $5,722$ connected components, with the largest component containing $73\%$ of the nodes, which amounts to $16,169$ nodes and $470,126$ edges. The second-largest component included only $0.04\%$ of the nodes. The NR network had a giant component, and this was the first indication of a small world network.
Next, to investigate if Wikidata was a small world network, we measured and analysed the average cluster coefficient and the average shortest path of NR. The average clustering coefficient was $0.6$ and the average shortest path was $3$. To compare the characteristics of the Wikidata network, we generated a small world network and a random network with a similar number of nodes and edges. In comparison, a similar small world network had similar characteristics with an average clustering coefficient of $0.1$ and an average shortest path of $3$. In contrast, a random network with similar characteristics had a much lower average clustering coefficient of only $0.004$ and an average shortest path of $3$. These comparisons, combined with the giant component, confirmed that NR formed a small world network. Information or resources could be disseminated quickly and efficiently across the network due to the short path lengths. Even though nodes were part of sub-groups, they could still reach other parts of the network with relatively few steps. Small-world networks are generally robust to random node removal. In the Wikidata case, this made it resilient to drop outs.
We then measured degree assortativity to investigate the tendency of editors to connect with editors with similar degree metric. NR network had assortativity at $-0.2$. This suggests a disassortative network, meaning that editors with high degree tended to connect with editors with low degree. This suggests an inclusive community with the flow of information or resources from well-connected nodes to the periphery.
Figure 8 presents the clustering coefficient versus the degree of a node in the NR network. The figure showed that a high number of nodes had both high clustering coefficient and degree. Such nodes were likely to be influential or central within their respective clusters or communities and lead to robust networks with multiple paths for information or interactions to flow. As a result, the network was resilient to the removal of individual nodes (i.e., editor drop out). As mentioned earlier, this could be the reason why NF, NH and NR presented similar clustering coefficient distributions in Section 6.1.1. Despite bots being the most connected editors in NF, their removal in NH and NR did not influence node connections.
We also measured the distribution of the shortest paths in the NR network. Similar to the MSN study [25], we randomly sampled $1000$ nodes and calculated their shortest paths to all other nodes (Figure 8). We observed that the longest shortest path was seven, which indicates that long paths did not exist in the NR network. By examining the percentile of shortest paths, we discovered that one can reach $90\%$ of nodes in just four steps. This indicates that the network was highly interconnected with short average path lengths. This reflects efficient communication, robustness, and the small-world nature of the network, where local clustering and global reach coexist.
@@
6.1.3 Investigation of registered editors characteristics
@@
Figure 9 illustrates the distribution of registered editors based on the number of edits, posts and account age. Our analysis revealed that both the number of edits and posts followed a power law distribution. Approximately $80\%$ of editors had fewer than $5$K edits and ten posts, suggesting low engagement with the community. Only a small percentage of editors ($20\%$) were actively participating in discussions, suggesting that they were potentially the ones making the key decisions within the Wikidata community.
@@
Furthermore, we analysed the distribution of account age and found that $24\%$ of editors had stayed in the community for less than a year, and $10\%$ of editors had remained active for more than ten years. This trend was consistent with the first four years of Wikidata [38].
We further determined which editors made high contributions to discussions based on their access level and account age (Figure 10). Editors who had been active on Wikidata for over six years were the most active in discussions. Property creators were the most active, followed by basic level users. This suggests that editors with basic level editing rights made a great share of contributions in discussions, such as helping others and participating in decisions, etc.
@@ 6.1.4 Summary
Furthermore, we found that only $20\%$ of editors were actively involved in discussions and decision-making, and $10\%$ of editors were members for more than six years. It is interesting to note that the majority of editors who participated in discussions were basic level users, and some of them had the highest number of posts.
6.2 RQ2 - What factors affect whether a discussion will receive a response?
The study aimed to investigate the factors that influence whether a post on item talk pages would receive a response. The ablation study included text embedding for the post content, graph embedding for the post topology (Gpost), and features for the editor who wrote the post (Table 4).
We conducted an experiment using different text embeddings to determine their effectiveness. We tested whether using text embeddings for only the first posts of threads (TfisrtPost) performed better than other posts within the threads (Tpost), or the entire thread from the first to the ith post (Tthread). If TfirstPost and Tpost showed different performances, it could indicate that the position of the post influences the prediction. Additionally, if Tthread and Tpost showed different performances, it could indicate that previous posts in the thread can affect the prediction.
The results of the experiment are presented in Table 5. The results showed a balance between precision and recall scores for all features, suggesting that the ML model had balance in correctly identifying positive instances and avoiding false positives. This means that the model was not heavily biased towards one type of error. The model was likely to generalise well to new data, suggesting that it was not overfitting to the training data. This was further supported by the F1 scores, in which a score similar to precision and recall indicates a balanced and high-performing model.
Table 5 demonstrates that, among individual features, graph embedding performed best, with $0.72\%$ accuracy, in predicting whether a post would receive a response. This result was expected since a post that was part of a highly connected thread had a higher chance of receiving a response. In addition, we found high performance, with approximately $0.7\%$ accuracy, for the text embeddings with minimal variations between the TfirstPost, Tpost, and Tthread. This suggested that the position of the posts and the previous posts did not have a significant influence, but the content of the text played a significant role in eliciting a response. The results mean that based on the content of the post, we could predict whether it receives a response. This was particularly important for TfirstPost results, which initiated a discussion.
The best features for predicting whether a post would receive a response were its content and where it was posted (Gpost + TfirstPost), with $0.79\%$ accuracy. The characteristics of the editor, such as their account age and access level, did not seem to significantly impact whether the post would receive a response. This suggests that it was more important to focus on the content of the post and where it was placed rather than who wrote it. This shows that the community was inclusive. These results are similar to previous findings about the communication characteristics of the registered editors in $RQ1$.
Table 5: Performance results of the ablation study for the features individually and in combination. The ML model predicted whether a post would receive a response.